A man is 24 years older than his son. Three years ago, he was eight years less than his son's age three years from now. How old is the man?
Answer by Boreal(15235)

son is now x
man's is x+24
three years ago, man was x+24-3=x+21
that is eight years less than his son's age three years from now, which will be x+3
so x+21 is 8 less than x+3
therefor x+21+8=x+3
so x+29 = x+3
this is not true so there is a mistake in the formulation of the problem.
Example: son is 6 man is 30
three years ago, son was 3 and man was 27
the son will be 9 3 years from now.